### v1

@@ 2 Preliminaries
Definition 6 ([14]).
A subset $C$ of a Hadamard manifold $\mathcal{M}$ is said to be a geodesically convex if, for any two points $x$ and $y$ in $C$, the geodesic joining $x$ to $y$ is contained in $C$, i.e., if $\gamma(x,y;\cdot)\colon[a,b]\to\mathcal{M}$ is a geodesic such that $x=\gamma(x,y;a)$ and $y=\gamma(x,y;b)$, then $\gamma((1-t)a+tb)\in C$ for all $t\in[0,1]$.
Definition 7 ([14]).
Let $\mathcal{M}$ be a Hadamard manifold. A function $f\colon\mathcal{M}\to\mathbb{R}$ is said to be a geodesically convex if, for any geodesic $\gamma(x,y;\cdot)\colon[a,b]\to\mathcal{M}$, the composition function $f\circ\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{R}$ is convex, i.e.,
@@
Proposition 8 ([44]).
@@
Remark 9.
Let $\mathcal{M}$ be a Hadamard manifold and $\triangle(x,y,z)$ be a geodesic triangle in $\mathcal{M}$. Then, we have the following:
  1. (i)
  2. (ii)
  3. (iii)
    $d^{2}(x,y)\leq\langle\exp_{x}^{-1}z,\exp_{x}^{-1}y\rangle+\langle\exp_{y}^{-1}z,\exp_{y}^{-1}x\rangle$.
Lemma 10 ([46]).
  1. (a)
  2. (b)
Lemma 11 ( [47]).
Let $\mathcal{M}$ be an Hadamard manifold and $x,y,z\in\mathcal{M}$. Then
$\|\exp_{x}^{-1}z-P_{x,z}\exp_{y}^{-1}z\|\leq d(x,y).$
Let $\mathcal{M}$ be a Hadamard manifold and let $m\in\mathbb{N}$. Then, the product space of manifold $\mathcal{M}$ is defined by
$\mathcal{M}^{N}=\{x=(x_{1},\ldots,x_{N})\colon x_{i}\in\mathcal{M}\text{ for }i=1,2,\ldots,N\}.$
Note that product space $\mathcal{M}^{N}$ is again a Hadamard manifold and the tangent space of $\mathcal{M}^{N}$ at $x=(x_{1},\ldots,x_{N})$ is $\mathcal{T}_{x}\mathcal{M}^{N}=\mathcal{T}_{x_{1}}\mathcal{M}\times\cdots\times\mathcal{T}_{x_{N}}\mathcal{M}$. The distance between $x=(x_{1},\ldots,x_{N}),y=(y_{1},\ldots,y_{N})\in\mathcal{M}^{N}$ is defined by
@@
Proposition 12 ([48]).
@@
The proximal operator $\operatorname*{Prox}_{\lambda f}$ is well defined for all $\lambda>0$.
Definition 15 ([4]).
Let $\mathcal{M}$ be a Hadamard manifold and let $T\colon\mathcal{M}\to\mathcal{M}$ be an operator. The operator $R_{T}\colon\mathcal{M}\to\mathcal{M}$ defined by
$R_{T}(x)=\exp_{T(x)}(-\exp_{T(x)}^{-1}(x))\text{ for all }x\in\mathcal{M}$
is called the reflection of the operator $T$.
@@
Theorem 16 ([4]).
@@
Proposition 17 ([51, 4]).
Let $C$ be a nonempty, closed and geodesically convex subset of Hadamard manifold $\mathcal{M}$ with constant curvature $\kappa$. Then
@@
Lemma 18 ([4]).
Let $\mathcal{M}$ be an Hadamard manifold, $\lambda>0$ and $c\in\mathcal{M}$. Define $g\colon\mathcal{M}\to\mathbb{R}$ by $g(x)=d(c,x)\text{ for all }x\in\mathcal{M}$. Then, $\operatorname*{Prox}_{\lambda g}(x)=\gamma(x,c,t)$, where $t=\min\big\{\tfrac{\lambda}{d(x,c)},1\big\}$.
Theorem 19 ([4]).
Let $\mathcal{M}$ be a symmetric Hadamard Manifold, $\lambda>0$ and $c\in\mathcal{M}$. Let $g\colon\mathcal{M}\to\mathbb{R}$ be defined by $g(x)=d(c,x)\text{ for all }x\in\mathcal{M}$. Then, the reflection $R_{\lambda g}$ is nonexpansive.
@@
For proof of Proposition 20, see Appendix A.
@@ 3 Efficient Douglas-Rachford Algorithms on Hadamard Manifolds
First, we introduce an method that incorporates an inertial extrapolation acceleration term, given as follows:
@@
For the convergence analysis of inertial Mann method (12), we assume that the sequence $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ satisfy the following conditions:
@@
Let $\mathcal{M}$ be a Hadamard manifold and let $T\colon\mathcal{M}\to\mathcal{M}$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence generated by inertial Mann method (12), where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). Then $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$.
@@
Let $v\in\operatorname*{Fix}(T)$. From (12) and Proposition 20, we have
@@
From (15) and (18), we have $d(x_{n},x_{n-1})=\|x_{n}^{\prime}-x_{n-1}^{\prime}\|$ and $d(x_{n},x_{n+1})=\|x_{n}^{\prime}-x_{n+1}^{\prime}\|$. Thus
@@
Therefore $\psi_{n}$ is nonincreasing.
Note $\psi_{1}=\phi_{1}-\theta_{1}\phi_{0}+K_{1}d^{2}(x_{1},x_{0})>0$. Since $\{\theta_{n}\}$ is bounded above by $\theta$, we get
@@
From (25) and (26), we obtain that
@@
$K\sum_{n=1}^{\infty}d^{2}(x_{n+1},x_{n})\leq\frac{\psi_{1}}{1-\theta},$ (27)
which implies that $\lim_{n\to\infty}d(x_{n},x_{n+1})=0$.
From (15) and (18), we have
@@
$=d(x_{n+1},x_{n})+\theta_{n}d(x_{n-1},x_{n})\to 0\text{ as }n\to\infty.$
Observe that
$d(x_{n},y_{n})\leq d(x_{n},x_{n+1})+d(x_{n+1},y_{n})\to 0\text{ as }n\to\infty.$
@@
From (27) and Lemma 3, we obtain that $\lim_{n\to\infty}d(x_{n},v)$ exists.
We may assume that $\lim_{n\to\infty}d(x_{n},v)=l>0$. It follows from $\lim_{n\to\infty}d(x_{n},y_{n})=0$, that $\lim_{n\to\infty}d(y_{n},v)=l.$ Hence from (22), we obtain that $\lim_{n\to\infty}d(y_{n},Ty_{n})=0.$ Hence $\{y_{n}\}$ satisfies the LEAF point property, therefore by Proposition 12, $\{y_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. Since $\lim_{n\to\infty}d(x_{n},y_{n})=0$, we have $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. ∎
@@
Theorem 22 is the extension of Theorem 3.3 of Sahu [29] from Hilbert spaces to the setting of Hadamard manifolds.
@@
Let $\mathcal{M}$ be a Hadamard manifold and let $T\colon\mathcal{M}\to\mathcal{M}$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence generated by inertial Mann method (12) with $x_{0}=x_{1}$, where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). Then $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$ with the following rate of convergence:
$R_{T,\{x_{n}\}}(n)\leq o\left(\frac{1}{\sqrt{n}}\right).$
@@
$\quad-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n}),$
which implies from condition (C1) that
@@
From (28), we have
@@
$[V_{n+1}]_{+}\leq\theta^{n}[V_{1}]+\sum_{j=1}^{n}\theta^{j-1}\delta_{n+1-j}.$ (33)
@@
By using the nonexpansivity of $T$, (12) and Proposition 13, we have
@@
$\sum_{i=1}^{n}d^{2}(x_{i+1},Tx_{i+1})\leq 5\sum_{i=1}^{n}d^{2}(y_{i},Ty_{i})\leq\frac{5}{a(1-b)}\left(1+\frac{\theta^{2}(1+\theta)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)}{K(1-\theta)}\right)d^{2}(x_{0},v).$
Thus
@@
Initialization: Choose $x_{0}$ arbitrary.
@@
$\begin{cases}y_{n}=\gamma(x_{n},Tx_{n};\alpha_{n}),\\ x_{n+1}=T^{p}(y_{n})\text{ for all }n\in\mathbb{N}_{0},\end{cases}$ (36)
@@
Note that this method is inspired from the $p$-accelerated normal S-iteration method (7); accordingly, we refer to it by the same name.
We now study the convergence of the $p$-accelerated normal S-iteration method (36).
@@
Let $C$ be a nonempty, closed and geodesically convex subset of a Hadamard manifold $\mathcal{M}$ and let $T\colon C\to C$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be the sequence generated by $p$-accelerated normal S-iteration method (36), where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$ with the convergence rate gives as follows:
@@
Let $v\in\operatorname*{Fix}(T)$. From (36) and Proposition 8, we have
@@
Hence, $\lim_{n\to\infty}d(x_{n},v)$ exists and $\{x_{n}\}$ is bounded.
Next we show that $\lim_{n\to\infty}d(x_{n},Tx_{n})=0$. From Proposition 13, we have
@@
$\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})\leq d^{2}(x_{n},v)-d^{2}(x_{n+1},v).$ (38)
Since $\lim_{n\to\infty}d(x_{n},v)$ exists, from (38) we have
$\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})<\infty,$ (39)
@@
$\liminf_{n\to\infty}d(x_{n},Tx_{n})=0,$ (40)
because otherwise $d(x_{n},Tx_{n})\geq K$ for all $n\in\mathbb{N}_{0}$ for some $K>0$ and then
which is a contradiction to the condition (39).
On the other hand, using the nonexpansivity of $T$ and Proposition 8, we have
$d(x_{n+1},Tx_{n+1})$ $=d(T^{p}(y_{n}),T^{p+1}(y_{n}))$ (41)
@@
$=(1-\alpha_{n})d(x_{n},Tx_{n})+\alpha_{n}d(Tx_{n},x_{n})=d(x_{n},Tx_{n}).$ (42)
This means that $\{d(x_{n},Tx_{n})\}$ is a monotone sequence. Combining this with (40) gives $\lim_{n\to\infty}d(x_{n},Tx_{n})=0$. Hence, $\{x_{n}\}$ satisfies the LEAF point property. Therefore, from Proposition 12, $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$.
Again from (38), we have
@@
From (42) sequence $\{d(x_{n},Tx_{n})\}$ is decreasing. It follows from (43) that
@@
Hence, $d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})}}\right).$
@@
If in addition, we take $0<a\leq\alpha_{n}\leq b<1$ in Theorem 27, then
$d(x_{n},Tx_{n})\leq\frac{d(x_{0},v)}{\sqrt{\sum_{i=1}^{n}a(1-b)}}=\frac{d^{2}(x_{0},v)}{\sqrt{na(1-b)}}.$
Thus $d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{n}}\right)$.
Remark 29.
While both Inertial Mann method (12) and $p$-accelerated normal S-iteration method (36) aim to compute a fixed point of the nonexpansive mapping $T$, their convergence behavior differs significantly in practice.
The $p$-accelerated normal S-iteration method (36) generates a relatively stable sequence, and the fixed-point residual $d(Tx_{n},x_{n})$ typically decreases in a monotonic fashion. As such, the value of $d(Tx_{n},x_{n})$ itself serves as a reliable indicator of the convergence rate.

### v2

@@ 2 Preliminaries
Hadamard manifolds have the nice feature that they resemble convexity properties from Euclidean spaces. A subset $C$ of a Hadamard manifold $\mathcal{M}$ is said to be geodesically convex if, for any two points $x$ and $y$ in $C$, the geodesic joining $x$ to $y$ is contained in $C$. A function $f\colon\mathcal{M}\to\mathbb{R}$ is said to be geodesically convex if, for any geodesic $\gamma(x,y;\cdot)\colon[a,b]\to\mathcal{M}$, the composition function $f\circ\gamma(x,y;\cdot)\colon[a,b]\to\mathbb{R}$ is convex, i.e.,
@@
The following proposition highlights the convexity of the distance function.
Proposition 7 ([Sakai]).
@@
Proposition 8 ([ShikherCNSNS2024]).
$d^{2}(u,z)\leq(1-t)d^{2}(x,z)+td^{2}(y,z)-t(1-t)d^{2}(x,y).$
Proposition 9.
(The proof is given in Appendix A.)
Let $\mathcal{M}$ be a Hadamard manifold and let $N\in\mathbb{N}$. Then, the product space of manifold $\mathcal{M}$ is defined by $\mathcal{M}^{N}=\{x=(x_{1},\ldots,x_{N})\colon x_{i}\in\mathcal{M}\text{ for }i=1,2,\ldots,N\}.$ Note that product space $\mathcal{M}^{N}$ is again a Hadamard manifold. The tangent space of $\mathcal{M}^{N}$ at $x=(x_{1},\ldots,x_{N})$ is $\mathcal{T}_{x}\mathcal{M}^{N}=\mathcal{T}_{x_{1}}\mathcal{M}\times\cdots\times\mathcal{T}_{x_{N}}\mathcal{M}$. The distance between two points $x=(x_{1},\ldots,x_{N})$ and $y=(y_{1},\ldots,y_{N})$ in $\mathcal{M}^{N}$ is defined by
@@
Moreover, the geodesic in $\mathcal{M}^{N}$ joining $x=(x_{1},\ldots,x_{N})$ to $y=(y_{1},\ldots,y_{N})$ is given componentwise by $\gamma(x,y;t)=\bigl(\gamma(x_{i},y_{i};t),\ldots,\gamma(x_{N},y_{N};t)\bigr)\text{ for all }t\in[0,1]$.
We now recall an important property related to the convergence of sequences satisfying the LEAF point (limit existence and approximate fixed point) property.
Proposition 10 ([ShikherCNSNS2024]).
@@
The proximal operator of a proper, geodesically convex and lower semicontinuous function is firmly nonexpansive operator (see [Jost1995, Martin3]), i.e., for any $x,y\in\mathcal{M}$ and $t\in[0,1]$
$d(\operatorname{Prox}_{\!\lambda f}(x),\operatorname{Prox}_{\!\lambda f}(y))\leq d(\gamma(x,\operatorname{Prox}_{\!\lambda f}(x);t),\gamma(y,\operatorname{Prox}_{\!\lambda f}(y);t)).$ (10)
@@
$R_{\lambda f}(x)=\exp_{\operatorname*{Prox}_{\lambda f}(x)}(-\exp_{\operatorname*{Prox}_{\lambda f}(x)}^{-1}(x))\text{ for all }x\in\mathcal{M}.$
That is, $R_{\lambda f}(x)$ is the reflection of $x$ with respect to $\operatorname*{Prox}_{\lambda f}(x)$ along the unique geodesic joining $x$ to $\operatorname*{Prox}_{\lambda f}(x)$.
Lemma 11 ([Bergmann2016]).
Let $\mathcal{M}$ be an Hadamard manifold, $\lambda>0$ and $c\in\mathcal{M}$. Define $f\colon\mathcal{M}\to\mathbb{R}$ by $f(x)=d(c,x)\text{ for all }x\in\mathcal{M}$. Then, $\operatorname*{Prox}_{\lambda f}(x)=\gamma(x,c,t)$, where $t=\min\big\{\tfrac{\lambda}{d(x,c)},1\big\}$.
[…]
Theorem 12 ([Bergmann2016]).
@@
Next, we present some results for functions with nonexpansive reflections.
Proposition 13 ([Leon2013, Bergmann2016]).
Let $C$ be a nonempty, closed and geodesically convex subset of Hadamard manifold $\mathcal{M}$ with constant sectional curvature. Then
@@
Theorem 14 ([Bergmann2016]).
Let $\mathcal{M}$ be a Hadamard Manifold, $\lambda>0$ and $c\in\mathcal{M}$. Let $g\colon\mathcal{M}\to\mathbb{R}$ be defined by $g(x)=d(c,x)\text{ for all }x\in\mathcal{M}$. Then, the reflection $R_{\lambda g}$ is nonexpansive.
@@
[…]
(The proof is given in Appendix B.)
@@ 3 Efficient Douglas-Rachford Algorithms on Hadamard Manifolds
First, we introduce a method that incorporates an inertial extrapolation acceleration term, given as follows:
@@
For the convergence analysis of inertial Mann method (11), we assume that the sequences $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ satisfy the following conditions:
@@
Remark 17.
The upper bound for the inertial parameter $\{\theta_{n}\}$ in the assumption of Boţ et al. [Bot2015] (see (9)) depends on several auxiliary parameters $\alpha$, $\delta$, and $\sigma$. Consequently, determining an admissible range for $\{\theta_{n}\}$ requires careful tuning of these coupled parameters and is not straightforward. In contrast, under our condition (C3), an explicit upper bound for $\{\theta_{n}\}$ can be directly obtained once an upper bound for the sequence $\{\alpha_{n}\}$ is fixed. Intuitively, the parameter $\theta_{n}$ controls the amount of inertia introduced at each iteration, while the bound involving $\epsilon=(1-b)/b$ ensures that this inertial effect remains sufficiently small relative to the relaxation parameter $\alpha_{n}$. This allows the parameters to be chosen independently, making the conditions easier to verify and more convenient to implement in practice than those in [Bot2015].
@@
Let $\mathcal{M}$ be a Hadamard manifold and let $T\colon\mathcal{M}\to\mathcal{M}$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence generated by inertial Mann method (11), where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). For $v\in\operatorname*{Fix}(T)$, define $\phi_{n}=d^{2}(x_{n},v),\psi_{n}=\phi_{n}-\theta_{n}\phi_{n-1}+K_{n}d^{2}(x_{n},x_{n-1})$, where $K_{n}=\theta_{n}(1+\theta_{n}+\epsilon(1-\theta_{n}))$. Assume that $\psi_{1}>0$. Then, we have the following:
  • (i)
    $\{\psi_{n}\}$ is nonincreasing.
  • (ii)
    $\sum_{n=1}^{\infty}d^{2}(x_{n+1},x_{n})\leq\frac{\psi_{1}}{K(1-\theta)},$ where $K=\epsilon-\theta(1+\epsilon+\max\{1,\epsilon\})>0$.
  • (iii)
    $\lim_{n\to\infty}d(x_{n},v)$ and $\lim_{n\to\infty}d(x_{n},v)$ exists and $\lim_{n\to\infty}d(y_{n},Ty_{n})=0$.
  • (iv)
    $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$.
@@
(i) Let $v\in\operatorname*{Fix}(T)$. From (11) and Proposition 9, we have
@@
using (14) and (17), we have $d(x_{n},x_{n-1})=\|x_{n}^{\prime}-x_{n-1}^{\prime}\|$ and $d(x_{n},x_{n+1})=\|x_{n}^{\prime}-x_{n+1}^{\prime}\|$. Thus
@@
Therefore $\{\psi_{n}\}$ is nonincreasing.
(ii) Note $\psi_{1}=\phi_{1}-\theta_{1}\phi_{0}+K_{1}d^{2}(x_{1},x_{0})>0$. Since $\{\theta_{n}\}$ is bounded above by $\theta$, we get
@@
It follows from (24) and (25) that
@@
$\sum_{n=1}^{\infty}d^{2}(x_{n+1},x_{n})\leq\frac{\psi_{1}}{K(1-\theta)}.$ (26)
(iii) It follows from (26) that
$\lim_{n\to\infty}d(x_{n},x_{n+1})=0.$ (27)
Using (14), (17) and (27), we deduce that
@@
$=d(x_{n+1},x_{n})+\theta_{n}d(x_{n-1},x_{n})\to 0\text{ as }n\to\infty.$ (28)
Observe from (27) and (28) that
$d(x_{n},y_{n})\leq d(x_{n},x_{n+1})+d(x_{n+1},y_{n})\to 0\text{ as }n\to\infty.$ (29)
@@
From (26) and Lemma 4, we obtain that $\lim_{n\to\infty}d(x_{n},v)$ exists. We may assume that $\lim_{n\to\infty}d(x_{n},v)=l>0$. It follows from (29) that $\lim_{n\to\infty}d(y_{n},v)=l.$ Hence from (13), we obtain that $\lim_{n\to\infty}d(y_{n},Ty_{n})=0.$
(iv) Note that $\lim_{n\to\infty}d(y_{n},v)=l$ exists and $\lim_{n\to\infty}d(y_{n},Ty_{n})=0$. Hence $\{y_{n}\}$ satisfies the LEAF point property, therefore by Proposition 10, $\{y_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. Since $\lim_{n\to\infty}d(x_{n},y_{n})=0$, we have $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. ∎
@@
Theorem 18 is the extension of Theorem 3.3 of Sahu [Sahusoft2020] from Hilbert spaces to the setting of Hadamard manifolds.
@@
Let $\mathcal{M}$ be a Hadamard manifold and let $T\colon\mathcal{M}\to\mathcal{M}$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be a sequence generated by inertial Mann method (11) with $x_{0}=x_{1}\in\mathcal{M}$, where $\{\alpha_{n}\}$ and $\{\theta_{n}\}$ are sequences satisfying the conditions (C1), (C2) and (C3). Then $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$ with the following rate of convergence:
$R_{T,\{x_{n}\}}(n)=o\left(\frac{1}{\sqrt{n}}\right).$
@@
$\quad-\alpha_{n}(1-\alpha_{n})d^{2}(y_{n},Ty_{n}).$
It follows from condition (C1) that
@@
$\leq-a(1-b)d^{2}(y_{n},Ty_{n})+\theta_{n}(1+\theta_{n})d^{2}(x_{n},x_{n-1}),$
which implies that
@@
$[V_{n+1}]_{+}\leq\theta^{n}[V_{1}]_{+}+\sum_{j=1}^{n}\theta^{j-1}\delta_{n+1-j}.$ (34)
@@
By using the nonexpansivity of $T$, (11) and Proposition 8, we deduce that
@@
$\sum_{i=1}^{n}d^{2}(x_{i+1},Tx_{i+1})\leq 5\sum_{i=1}^{n}d^{2}(y_{i},Ty_{i})\leq\frac{5}{a(1-b)}\left(1+\frac{\theta^{2}(1+\theta)}{K(1-\theta)^{2}}+\frac{\theta(1+\theta)}{K(1-\theta)}\right)d^{2}(x_{0},v),$ (37)
which implies that $\sum_{i=1}^{n}(R_{T,\{x_{n}\}}(n))^{2}<\infty$. It follows from Lemma 1 that
$R_{T,\{x_{n}\}}(n)=o\left(\frac{1}{\sqrt{n}}\right).$
Moreover, by (37), we obtain
@@
Initialization: Choose $x_{1}\in C$ arbitrary.
@@
$\begin{cases}y_{n}=\gamma(x_{n},Tx_{n};\alpha_{n}),\\ x_{n+1}=T^{p}(y_{n})\text{ for all }n\in\mathbb{N},\end{cases}$ (38)
@@
Note that this method is inspired from the $p$-accelerated normal S-iteration method (7); accordingly, we refer to it by the same name. We now study the convergence of the $p$-accelerated normal S-iteration method (38).
@@
Let $C$ be a nonempty, closed and geodesically convex subset of a Hadamard manifold $\mathcal{M}$ and let $T\colon C\to C$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be the sequence generated by $p$-accelerated normal S-iteration method (38), where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then, we have the following:
  1. (i)
    $\lim_{n\to\infty}d(x_{n},v)$ exists for any $v\in\operatorname*{Fix}(T)$.
  2. (ii)
    $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})<\infty$.
  3. (iii)
    $\{d(x_{n},Tx_{n})\}$ is decreasing.
  4. (iv)
    $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$.
@@
(i) Let $v\in\operatorname*{Fix}(T)$. From (38) and Proposition 7, we have
@@
(ii) From Proposition 8, we have
@@
$\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})\leq d^{2}(x_{n},v)-d^{2}(x_{n+1},v).$
Summing up the above inequality from $n=1$ to $k$, we have
$\sum_{n=1}^{k}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})\leq\sum_{n=1}^{k}(d^{2}(x_{n},v)-d^{2}(x_{n+1},v))=d^{2}(x_{0},v),$ (40)
@@
$\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})<\infty.$ (41)
(iii) Using the nonexpansivity of $T$ and Proposition 7, we have
$d(x_{n+1},Tx_{n+1})$ $=d(T^{p}(y_{n}),T^{p+1}(y_{n}))$
@@
$=(1-\alpha_{n})d(x_{n},Tx_{n})+\alpha_{n}d(Tx_{n},x_{n})=d(x_{n},Tx_{n}).$
(iv) Using (41), we have
$\liminf_{n\to\infty}d(x_{n},Tx_{n})=0,$ (42)
because otherwise $d(x_{n},Tx_{n})\geq K$ for all $n\in\mathbb{N}$ for some $K>0$ and then
which is a contradiction to the condition (41). From (iii) sequence $\{d(x_{n},Tx_{n})\}$ is decreasing. Combining this fact with (42) gives $\lim_{n\to\infty}d(x_{n},Tx_{n})=0$. Hence, $\{x_{n}\}$ satisfies the LEAF point property. Therefore, from Proposition 10, $\{x_{n}\}$ converges to an element of $\operatorname*{Fix}(T)$. ∎
Theorem 24.
Let $C$ be a nonempty, closed and geodesically convex subset of a Hadamard manifold $\mathcal{M}$ and let $T\colon C\to C$ be a nonexpansive operator such that $\operatorname*{Fix}(T)\neq\emptyset$. Let $\{x_{n}\}$ be the sequence generated by $p$-accelerated normal S-iteration method (38), where $\{\alpha_{n}\}$ is a sequence in $(0,1)$ satisfying the condition $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$. Then, we have the following:
  1. (i)
    $d(x_{n},Tx_{n})\leq\frac{d(x_{0},v)}{\sqrt{\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})}}$ and $d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{\sum_{i=1}^{n}\alpha_{i}(1-\alpha_{i})}}\right)$.
  2. (ii)
    If in addition $0<a\leq\alpha_{n}\leq b<1$, then $d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{n}}\right)$.
Proof.
(i) Using (40), we have
@@
From Theorem 23(iii), sequence $\{d(x_{n},Tx_{n})\}$ is decreasing. It follows from (43) that
@@
From Theorem 23(ii), we have $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})d^{2}(x_{n},Tx_{n})<\infty$. Since $\sum_{n=1}^{\infty}\alpha_{n}(1-\alpha_{n})=\infty$, it follows from Lemma 1 that
(ii) Since $0<a\leq\alpha_{n}\leq b<1$, from Theorem 23(ii), we have $\sum_{n=1}^{\infty}a(1-b)d^{2}(x_{n},Tx_{n})<\infty$ or $\sum_{n=1}^{\infty}d^{2}(x_{n},Tx_{n})<\infty$. Again it follows from Lemma 1 that
$d(x_{n},Tx_{n})=o\left(\frac{1}{\sqrt{n}}\right).$
@@
While both Inertial Mann method (11) and $p$-accelerated normal S-iteration method (38) aim to compute a fixed point of the nonexpansive mapping $T$, their convergence behavior differs significantly in practice. The $p$-accelerated normal S-iteration method (38) generates a relatively stable sequence, and the fixed-point residual $d(Tx_{n},x_{n})$ typically decreases in a monotonic fashion. As such, the value of $d(Tx_{n},x_{n})$ itself serves as a reliable indicator of the convergence rate.